Most of the crew retired a little earlier last night, after quite an eventful
The plan was to have a slightly less ambitious schedule today to give the crew a
chance to take stock and catch up with individual tasks.
Dario helped me clean and sort the contents of the EVA prep room, Christiane and
Carmel worked on producing an inventory of the engineering area, Cyprien worked
on a report for ... and Vince worked in the Green hab. The HSO and I also spent
a bit of time looking at contingency measures that we could take if there is a
deterioration in the toilet functionality.
In the mid-morning we paused for demonstration/practice of injury evacuation
procedures, with Carmel demonstrating how to improvise a stretcher using three
pack harnesses, and then several alternative methods of moving an injured crew
member were discussed and practised.
All the while we were waiting for news from Earth regarding the arrival of our
long anticipated replacement parts for our space suits
EVA3 departed the hab as planned, except I decided to substitute Carmel for
Cyprien. The substitution was because I wanted Carmel's experience with ATVs to
help with the checking of oil and investigation of the problems experienced the
previous afternoon. I had also been informed by Cyprien that he had decided that
he would prefer to do the remaining sampling further from the hab.
The EVA departed the hab two at a time so as to allow the solar cooker to be
accompanied in the airlock. The first two EVA members exited the airlock at
14:05, and the final two exited at 14:23.
Christiane and Carmel climbed the slope at the rear of the hab and immediately
started to assemble the solar cooker.
Dario deployed his drone to start practising his flight skills. When he had
finished this, I asked him to assist Christiane so that Carmel could look at the
ATVs with me.
I found that I was able to start ATV3 (which was the one towed back to the hab
yesterday) without any problem, and was able to drive from the hab to the road
and back without any problems.
ATV4, on the other hand, which is the one I had experienced problems with,
continued to misbehave. I found I could start it, but after driving a short
distance it would lose power unless applied much more throttle than I wanted. So
my choices were too fast or stopped. Carmel then tried the same machine, and
initially had no drive at all, much as I experienced at departing our first
waypoint yesterday. A more thorough description will be included in the
I had asked Habcom to take the opportunity to test a seventh radio that I had
found when cleaning the EVA prep room, and our tests revealed that it worked for
periods, but occasionally stopped working (with the display going blank)and
required a turn-off followed by turn-on to resume operation. I asked him to
attach a tag indicating this behaviour.
The oil and fuel was checked on all ATV's, and the oil was a little low in ATV1
(25% full), but we found no problem with either of the ATVs which had given
At around 16:00 Dario came running down from the solar cooker exclaiming that
there was a medical emergency. I heard him say that it was not a drill, and
Carmel and I ran up to assess the situation while Dario spoke to the hab. We had
begun tending to the stricken party and I had requested immediate attendance by
the HSO when the I was informed that this was in fact a drill and that
Christiane was following the HSOs earlier directive for members of the crew to
initiate drills without his knowledge.
At that point we resumed sim, and proceeded with the exercise in sim, escorting
a heat stroke effected crewman back to the airlock and then carrying out
appropriate first aid procedures.
At 16:34 the EVA re-exited the airlock and completed their tasks. A6 16:55 the
solar cooker team announced that they had produced pot of near boiling water
which could be used for preparing dinner.
All that remained was to complete our regular external engineering duties, and
at 17:30, the EVA returned to the airlock.
We had not planned anything particularly ambitious for today, but it still felt
like there were not enough hours in the day, and I think we will all sleep well
Supplemental Report on Helmets and Radios
Four refurbished helmets have been received. All have been tested and
all are working well, with good visibility.
Two of the four neck rings have have breaks which have been repaired
with tape, which has resulted in quite a tight fit. There were
insufficient small straps of the type used to attach neck rings to
back packs, so we had to improvise with some different straps.
Several pictures from todays reports illustrate the end result.
Six radios with headsets were also received. The radios have been
tested and work well over short distances. We have not yet been able
to test the range of the radios as they were received after the end of
The headsets work, although the volume seems to be greatly attenuated
when in use, and the ear fastenings to not look sufficiently secure to
be used without additional support. They are quite difficult to put
on. But they work and we will be happy to use them.